# DisCoPatch: Batch Statistics Are All You Need For OOD Detection, But Only If You Can Trust Them

### v1

@@ Abstract
Out-of-distribution (OOD) detection holds significant importance across many applications. While semantic and domain-shift OOD problems are well-studied, this work focuses on covariate shifts - subtle variations in the data distribution that can degrade machine learning performance. We hypothesize that detecting these subtle shifts can improve our understanding of in-distribution boundaries, ultimately improving OOD detection. In adversarial discriminators trained with Batch Normalization (BN), real and adversarial samples form distinct domains with unique batch statistics — a property we exploit for OOD detection. We introduce DisCoPatch, an unsupervised Adversarial Variational Autoencoder (VAE) framework that harnesses this mechanism. During inference, batches consist of patches from the same image, ensuring a consistent data distribution that allows the model to rely on batch statistics. DisCoPatch uses the VAE’s suboptimal outputs (generated and reconstructed) as negative samples to train the discriminator, thereby improving its ability to delineate the boundary between in-distribution samples and covariate shifts. By tightening this boundary, DisCoPatch achieves state-of-the-art results in public OOD detection benchmarks. The proposed model not only excels in detecting covariate shifts, achieving 95.5% AUROC on ImageNet-1K(-C), but also outperforms all prior methods on public Near-OOD (95.0%) benchmarks. With a compact model size of $\leq$ 25MB, it achieves high OOD detection performance at notably lower latency than existing methods, making it an efficient and practical solution for real-world OOD detection applications. The code will be made publicly available.
@@ 3.1 Overview
A prevalent generative-based approach for OOD detection involves utilizing the trained generator to evaluate the likelihood of unseen samples. However, in adversarial setups, some information about the ID boundary will be incorporated into the discriminator, as it learns to assess the probability of a sample being real (ID) or synthetic (OOD). As mentioned in Section 1, in this paper, we exploit the observation that BN can help an adversarially trained discriminator to separate underlying data distributions by recognizing that clean and adversarial images are drawn from two distinct domains (i.e. ID and OOD), in such way that that it can provide a boundary for the ID set. By adjusting where the discriminator learns to draw this boundary, we can create an OOD detector.
It is on this premise that we propose a Discriminative Covariate Shift Patch-based Network, DisCoPatch. DisCoPatch is an Adversarial VAE-inspired architecture, as shown in Figure 2, in which both the VAE and the discriminator are trained adversarially. DisCoPatch’s approach combines generative and reconstruction-based strategies to distill information about the ID set and OOD boundaries to the discriminator during training in an unsupervised manner. Unlike traditional adversarial methods, DisCoPatch’s focus is on leveraging the generator’s output as a tool to refine the discriminator. DisCoPatch’s discriminator only utilizes the current batch’s (of patches) statistics in the BatchNorm2D layer (i.e. parameter $m$ in Eq. (2) is set to one). Subsection 4.3 details extensive ablation experiments to demonstrate its effectiveness.
@@ 3.3 Patching Strategy
The patching strategy begins by taking a high-resolution input image, typically a standard 256$\times$256 resolution, and cropping it into $N$ random patches, each of size 64$\times$64. This approach allows the model to capture fine-grained details in different image regions. During training, batches are composed of patches sourced from multiple images rather than from a single one. This setup accelerates training and ensures that the model learns consistent ID features across a range of images, minimizing the potential for overfitting on individual image characteristics. During inference, however, each batch is constructed from $N$ patches taken from the same image to achieve independent results per image. The final anomaly score for an image is the mean of the scores of all patches within that same image. For the remainder of the paper, we refer to the model as DisCoPatch-$N$, indicating the number of patches per image used during inference. Both strategies are illustrated in Figure 4.
Figure 4: Different patching strategies employed by the model during training and inference. Only patches from a single image are used to determine the anomaly score during inference.
@@ 4.2 Evaluation Metrics
The evaluation metrics employed in OpenOOD by [64] are adopted for this work. These two main evaluation metrics are: (1) AUROC, which measures the area under the Receiver Operating Characteristic (ROC) curve, and displays the relationship between True Positive Rate (TPR) and False Positive Rate (FPR); and (2) FPR95, which measures the FPR when the TPR is equal to 95%, with lower scores indicating better performance. The full results are provided in the form "AUROC/FPR95%".
4.3 Batch Normalization Bias Analysis
We conduct various experiments to investigate the effect of Batch Normalization’s reliance on the batch statistics $\mu_{B}$ and $\sigma_{B}$ (Eq. 2) during training. We conduct the experiments using the complete ImageNet-1K dataset, center-cropped and resized to $256\times 256$ pixels, within the DisCoPatch framework. Unlike our main experiments that rely on patch-based batches, here we evaluate the model’s sensitivity to batch-level statistics across the full-resolution images. This model was used with the standard BatchNorm2D parameters, with track_running_stats set to True and the default momentum of 0.1.
In the first experiment, the model is analyzed in evaluation mode, as described in Subsection 2.6. In this mode, the model utilizes the running mean, $\mu_{R}$, and variance, $\sigma_{R}$, learned during training and neglects the statistics of the current batch. As mentioned in Section 1, this paper hypothesizes that the batch statistics of the batch being evaluated can be a powerful feature for OOD detection. Consequently, we also introduce an additional model, in which the track_running_stats option in PyTorch’s BatchNorm2D layer is set to False. This setting causes the model to disregard the running mean and variance obtained during training, and instead, it employs solely the statistics $\mu_{B}$ and $\sigma_{B}$ of the batch being tested for normalization. It can be noted that this setting does not update any of the layer weights, serving only as a tool to demonstrate the reliance on batch-level statistics during inference.
@@ 4.4 Baseline Models
For the ImageNet-1K benchmark, we compare our method, DisCoPatch, against SOTA public models, such as MOODv2 [30], NNGuide [39], and SCALE [62], all of which have demonstrated SOTA performance [^1] OpenOOD Benchmark https://zjysteven.github.io/OpenOOD/ on Near-OOD and Far-OOD detection for these datasets. These models employ feature-based and logit-based strategies to perform OOD detection. In the case of NNGuide, we evaluate three of its available backbones to assess performance across different model types: the top-performing RegNet [45], a more efficient ResNet-50 [15], and the lightweight MobileNetV2 [49]. The implementation details are provided in Appendix B.
@@ 4.5 Experimental Significance
5 Results
5.1 Batch Normalization Bias
Mode Near-OOD Far-OOD
Learned Statistics 38.4/98.0 34.5/98.1
Batch Statistics BS=1 64.9/87.4 70.9/71.2
BS=16 90.2/55.5 91.1/40.0
BS=32 95.7/28.6 93.4/36.2
BS=64 99.3/2.2 96.1/23.8
BS=128 99.8/0.3 97.4/17.0
BS=256 100.0/0.0 98.1/12.5
5.2 ImageNet-1K Benchmark
Table 2 shows that DisCoPatch-64 surpasses state-of-the-art methods in Near-OOD and Covariate Shift benchmarks, with a particularly large performance gap in the Covariate Shift OOD detection task. Although DisCoPatch does not achieve SOTA performance in any of the Far-OOD benchmarks, it is close to matching the best performers on the iNaturalist and DTD tasks and attains competitive performance in OpenImage-O, whilst being a much smaller model. Detailed performance analysis for Covariate Shift can be found in Appendix D.
@@ 5.2 ImageNet-1K Benchmark
MOODv2 [30] *85.0/58.1 *92.7/38.2 99.6/1.8 94.3/24.7 97.4/13.6 *70.5/73.9
SCALE [62] 77.4/67.7 85.4/51.8 98.0/9.5 97.6/11.9 94.0/28.2 *83.3/54.1
@@
DisCoPatch-64 (Prop.) 95.8/19.8 94.3/39.0 99.1/3.6 96.4/18.9 94.4/29.7 97.2/10.6
@@
The superiority of DisCoPatch in OOD detection extends beyond raw detection performance. As illustrated in Figure 5, DisCoPatch outperforms all other models in Near-OOD detection performance while reducing latency by as much as one order of magnitude. Even with only 16 patches per image, DisCoPatch outperforms all other models tested. However, increasing the patch count beyond 64 did not result in any noticeable performance improvements. The performance gap between DisCoPatch and the second fastest model, NNGuide (MobileNetV2 backbone), is near 25%.
@@ 7 Limitations & Future Work
Comprehensive benchmarking of Covariate Shift detection across diverse architectures is vital to advance the field. While DisCoPatch shows promise for generative-based setups, evaluating more models will clarify how architecture impacts detection performance. Covariate Shift detection is particularly critical in fields like medical imaging, where identifying distribution shifts is critical for ensuring trustworthy predictions. Expanding such evaluations to high-stakes domains could expand the relevance of DisCoPatch and similar models for robust, real-world applications.
While current results demonstrate the effective use of batch statistics for OOD detection, a signal-processing-focused analysis of feature propagation and suppression at each layer could provide key insights. This could reveal patterns specific to different shifts, improving model interpretability. We also aim to test more complex discriminator architectures to further explore these dynamics.
@@ 8 Conclusion
The supplementary material is organized as follows: Appendix A covers the datasets used in this work. Appendix B describes the implementation details of the employed models while Appendix C covers the compute resources required for training and evaluating the models. Appendix D provides detailed results on the ImageNet-1K Covariate Shift OOD detection benchmark.
@@ Appendix B Implementation Details
This appendix provides details of the implementation of the models employed in the paper for better reproducibility.
B.1 MOODv2
For the implementation of MOODv2, we follow the guidelines provided in[30]. Moreover, we have employed their official repository [^5] https://github.com/dvlab-research/MOOD under no official license. This work reduces the complexity of MOODv1 [29] while increasing the detection performance. In MOODv1, three steps are required. First, the Masked Image Modeling Vision Transformer (ViT) is pretrained on the ImageNet-21k [47] dataset. The second step involves fine-tuning the ViT on the same dataset. The third and final step consists of fine-tuning the ViT on the ID dataset. It can be observed that this process is costly when dealing with a substantial number of ID datasets. However, through experimental validation, MOODv2 has demonstrated that a well-prepared masked image modeling model does not require additional fine-tuning.
The selected encoder for MOODv2 is a BEiTv2 [40], which is already pre-trained and fine-tuned on ImageNet-21k, as provided in the aforementioned repository. Regarding the OOD score function, following the author’s recommendations, ViM [55] is utilized, which merges features and logits extracted from the trained image encoder. Here, $l_{i}$ represents the $i$-th logit of feature $x$ in the training set $X$; $\alpha$ denotes a model-specific constant; $R$ (with dimensions $N\times(N-D)$) corresponds to the portion of the eigenvector matrix $Q$ of $X$, ranging from the $(D+1)$-th column to the last, where $N$ stands for the principal dimension; and $C$ signifies the number of classes. Mathematically, the score can be expressed as
$s(x)=\frac{e^{\alpha\sqrt{x^{T}RR^{T}x}}}{\sum_{i=1}^{C}e^{l_{i}}+e^{\alpha\sqrt{x^{T}RR^{T}x}}}.$ (7)
B.2 SCALE
When employing the SCALE framework, we have followed the approach described in [62]. Moreover, we have used with the code made publicly available  [^6] https://github.com/kai422/SCALE (this repository is provided under the MIT License). It should be noted that SCALE is a post-hoc enhancement method for OOD detection which focuses on scaling network activations, rather than pruning them. The scaling factor applied to the activations $r$ is applied uniformly across all features, preserving the model’s logit ordinality and maintaining ID accuracy. Mathematically, the calculation of the logits with the scaled activations can be formulated as
$z^{\prime}=W\cdot(a\circ sf(a))+b,\text{ where }sf(a)_{j}=\exp(r).$ (8)
Furthermore, SCALE incorporates a training-time enhancement technique called Intermediate Tensor Shaping (ISH). ISH applies the same scaling concept during training to enhance OOD detection by emphasizing samples with stronger ID characteristics.
B.3 NNGuide
For the implementation of NNGuide, we have followed the guidelines in [39]. Moreover, we also have employed the publicly available code  [^7] https://github.com/roomo7time/nnguide, which is released under the Apache 2.0 License.
The technique NNGuide is a post-hoc, training-free inference method designed to improve classifier-based OOD detection scores by leveraging nearest neighbors in the ID dataset. This method aims to mitigate the overconfidence issue in Far-OOD samples while preserving fine-grained detection for Near-OOD instances. This is achieved by augmenting a classifier’s confidence score, $S_{\text{base}}(x)$, by using a guidance term $G(x)$, which is the confidence-weighted average similarity of the nearest neighbors. This term ensures that the score respects the data manifold’s boundary geometry. The guidance term is defined as the average similarity between the test input and its high-confidence nearest neighbors and can be formulated as
$S_{\text{NNGuide}}(x)=S_{\text{base}}(x)\cdot G(x).$ (9)
B.4 DisCoPatch
Table 3: Hyperparameters used for DisCoPatch’s training.
Model Lat. Dim. Hidden Dimensions lr
Full-Size 1024 32, 64, 128, 256, 512, 1024 $5e^{-5}$
Patches 1024 128, 256, 512, 1024 $8.5e^{-5}$
@@ C.2 Inference
MOODv2 86,530,984 19.26
SCALE 25,557,032 11.27
@@
Appendix D Covariate Shift Results on ImageNet-1K

### v3

# DisCoPatch: Taming Adversarially-driven Batch Statistics for Improved Out-of-Distribution Detection

@@ Abstract
Out-of-distribution (OOD) detection holds significant importance across many applications. While semantic and domain-shift OOD problems are well-studied, this work focuses on covariate shifts - subtle variations in the data distribution that can degrade machine learning performance. We hypothesize that detecting these subtle shifts can improve our understanding of in-distribution boundaries, ultimately improving OOD detection. In adversarial discriminators trained with Batch Normalization (BN), real and adversarial samples form distinct domains with unique batch statistics — a property we exploit for OOD detection. We introduce DisCoPatch, an unsupervised Adversarial Variational Autoencoder (VAE) framework that harnesses this mechanism. During inference, batches consist of patches from the same image, ensuring a consistent data distribution that allows the model to rely on batch statistics. DisCoPatch uses the VAE’s suboptimal outputs (generated and reconstructed) as negative samples to train the discriminator, thereby improving its ability to delineate the boundary between in-distribution samples and covariate shifts. By tightening this boundary, DisCoPatch achieves state-of-the-art results in public OOD detection benchmarks. The proposed model not only excels in detecting covariate shifts, achieving 95.5% AUROC on ImageNet-1K(-C), but also outperforms all prior methods on public Near-OOD (95.0%) benchmarks. With a compact model size of $\leq$ 25MB, it achieves high OOD detection performance at notably lower latency than existing methods, making it an efficient and practical solution for real-world OOD detection applications. The code is available at github.com/caetas/DisCoPatch.
@@ 3.1 Overview
A prevalent generative-based approach for OOD detection involves utilizing the trained generator to evaluate the likelihood of unseen samples. However, in adversarial setups, some information about the ID boundary will be incorporated into the discriminator, as it learns to assess the probability of a sample being real (ID) or synthetic (OOD). As mentioned in Section 1, we exploit the observation that BN can help an adversarially trained discriminator to separate underlying data distributions by recognizing that clean and adversarial images are drawn from two distinct domains (i.e. ID and OOD). Instead of solely estimating likelihoods, DisCoPatch clusters ID features while excluding OOD samples by employing an Adversarial VAE that generates and reconstructs samples to enhance feature separation. The discriminator penalizes poorly generated and reconstructed images, reinforcing the distinction between ID and OOD and refining the decision boundary.
It is on this premise that we propose a Discriminative Covariate Shift Patch-based Network, DisCoPatch. DisCoPatch is an Adversarial VAE-inspired architecture, as shown in Figure 2, in which both the VAE and the discriminator are trained adversarially. Adversarially trained discriminators have been previously used for OOD detection [27]. DisCoPatch’s approach combines generative and reconstruction-based strategies to distill information about the ID set and OOD boundaries to the discriminator during training in an unsupervised manner. Unlike traditional adversarial methods, DisCoPatch’s focus is on leveraging the generator’s output as a tool to refine the discriminator. DisCoPatch’s discriminator only utilizes the current batch’s (of patches) statistics in the BatchNorm2D layer (i.e. parameter $m$ in Eq. (2) is set to one). Subsection 4.6 details extensive ablation experiments to demonstrate its effectiveness.
@@ 3.3 Patching Strategy
The patching strategy begins by taking a high-resolution input image, typically a standard 256$\times$256 resolution, and cropping it into $N$ random patches, each of size 64$\times$64. This approach allows the model to capture fine-grained details in different image regions. During training, batches are composed of patches sourced from multiple images rather than from a single one. This setup accelerates training and ensures that the model learns consistent and robust ID features across a range of images, minimizing the potential for overfitting on individual image characteristics.
In this setup, using the standard BatchNorm2D layer constrains batch processing during inference to a single image, because normalization must be applied only to features from the same distribution; mixing ID and OOD samples in a batch would distort the statistics and hinder effective separation. To address this, we introduce a custom layer, PatchNorm2D, which retains the weights and biases of BatchNorm2D but enables the normalization of multiple groups of patches simultaneously. Specifically, each batch consists of several groups of $N$ patches, where each group contains patches from the same image that are normalized together. This ensures that the model produces independent results per image while maintaining the benefits of batch normalization. The final anomaly score for an image is the mean of the scores of all patches within that same image. For the remainder of the paper, we refer to the model as DisCoPatch-$N$, indicating the number of patches per image used during inference. Both strategies are illustrated in Figure 4.
Figure 4: Different patching strategies employed by the model. Training normalizes patches from multiple images, while inference normalizes patches per image.
@@ 4.2 Evaluation Metrics
The evaluation metrics employed in OpenOOD by [70] are adopted for this work. These two main evaluation metrics are: (1) AUROC, which measures the area under the Receiver Operating Characteristic (ROC) curve, and displays the relationship between True Positive Rate (TPR) and False Positive Rate (FPR); and (2) FPR95, which measures the FPR when the TPR is equal to 95%, with lower scores indicating better performance. The full results are provided in the form "AUROC/FPR95%". In addition to these quantitative metrics, a UMAP visualization of the extracted features is included as a qualitative evaluation, providing insight into the separability of ID and OOD samples in the learned feature space.
@@ 4.3 Baseline Models
For the ImageNet-1K benchmark, we compare our method, DisCoPatch, against SOTA public models, such as MOODv2 [34], NNGuide [44], SCALE [68], RankFeat&Weight [57], ASH [7], and FDBD [29], all of which have demonstrated SOTA performance [^1] OpenOOD Benchmark https://zjysteven.github.io/OpenOOD/ on Near-OOD and Far-OOD detection for these datasets. These models employ feature-based and logit-based strategies to perform OOD detection. In the case of NNGuide, we evaluate three of its available backbones to assess performance across different model types: the top-performing RegNet [49], a more efficient ResNet-50 [17], and the lightweight MobileNetV2 [53].
4.4 DisCoPatch Implementation & Compute
Appendix B contains the implementation specifics of DisCoPatch. The computational resources for training and evaluating models are detailed in Appendix C.
@@ 4.5 Experimental Significance
4.6 Batch Normalization Bias Analysis
We conduct various supplemental experiments to investigate the effect of Batch Normalization’s reliance on the batch statistics $\mu_{B}$ and $\sigma_{B}$ (Eq. 2) during training. We conduct the experiments using the complete ImageNet-1K dataset, center-cropped and resized to $256\times 256$ pixels, within the DisCoPatch framework. Unlike our main experiments that rely on patch-based batches, here we evaluate the model’s sensitivity to batch-level statistics across the full-resolution images. This model was used with the standard BatchNorm2D parameters, with track_running_stats set to True and the default momentum of 0.1.
In the first experiment, the model is analyzed in evaluation mode, as described in Subsection 2.6. In this mode, the model utilizes the running mean, $\mu_{R}$, and variance, $\sigma_{R}$, learned during training and neglects the statistics of the current batch. As mentioned in Section 1, this paper hypothesizes that the batch statistics of the batch being evaluated can be a powerful feature for OOD detection. Consequently, we also introduce an additional model, in which the track_running_stats option in PyTorch’s BatchNorm2D layer is set to False. This setting causes the model to disregard the running mean and variance obtained during training, and instead, it employs solely the statistics $\mu_{B}$ and $\sigma_{B}$ of the batch being tested for normalization. This configuration does not modify the layer weights and is solely used to illustrate dependency on batch-level statistics during inference. For a comprehensive analysis, refer to Appendix D. Alternative normalization strategies are covered in Appendix E.
@@ 4.6 Batch Normalization Bias Analysis
MOODv2 [34] (BEiTv2) *85.0/58.1 *92.7/38.2 99.6/1.8 94.3/24.7 97.4/13.6 *70.5/73.9
SCALE [68] (ResNet-50) 77.4/67.7 85.4/51.8 98.0/9.5 97.6/11.9 94.0/28.2 *83.3/54.1
@@
RankFeat [56] (ResNetv2-101) *89.4/47.9 *90.0/39.3 96.0/13.0 95.0/25.4 *92.4/33.3 *91.9/36.7
ASH [10] (ResNet-50) *74.1/80.6 *83.0/64.2 97.9/11.5 97.6/11.9 *92.8/32.7 *84.7/51.3
FDBD [37] (ResNet-50) *68.4/83.4 *81.4/66.1 97.6/12.4 98.0/10.7 *91.7/35.6 *82.2/57.8
DisCoPatch-64 (Proposed) 95.8/19.8 94.3/39.0 99.1/3.6 96.4/18.9 94.4/29.7 97.2/10.6
5 Results
5.1 ImageNet-1K Benchmark
Table 1 shows that DisCoPatch-64 surpasses state-of-the-art methods in Near-OOD and Covariate Shift benchmarks, with a particularly large performance gap in the Covariate Shift OOD detection task. Although DisCoPatch does not achieve SOTA performance in any of the Far-OOD benchmarks, it is close to matching the best performers on the iNaturalist and DTD tasks and attains competitive performance in OpenImage-O, whilst being a much smaller model. Detailed performance analysis for Covariate Shift can be found in Appendix G.
@@ 5.1 ImageNet-1K Benchmark
The superiority of DisCoPatch in OOD detection extends beyond raw detection performance. As illustrated in Figure 5, DisCoPatch outperforms all other models in Near-OOD detection performance while reducing latency by as much as one order of magnitude. Even with only 16 patches per image, DisCoPatch outperforms all other models tested. Appendix F offers a detailed examination of the performance of additional patch counts. However, increasing the patch count beyond 64 did not result in any noticeable performance improvements. The performance gap between DisCoPatch and the second fastest model, NNGuide (MobileNetV2 backbone), is near 25%.
@@
Figure 7 presents a UMAP visualization of the average features extracted by the discriminator from 64 patches per image. The results show that samples from Near-OOD datasets (NINCO, SSB-hard) cluster closer to the ID set (ImageNet-1K), while Far-OOD samples (OpenImage-O, DTD, iNaturalist) are mapped farther away. While the proposed method effectively structures the feature space into distinct clusters, the decision boundary remains suboptimal. This highlights the potential for integrating post-hoc feature-based methods to further enhance OOD separation.
Figure 7: UMAP visualization of discriminator-extracted features, each point corresponding to the mean features of image patches.
@@ 7 Limitations & Future Work
Comprehensive benchmarking of Covariate Shift detection across diverse architectures is vital to advance the field. While DisCoPatch shows promise for generative-based setups, evaluating more models will clarify how architecture impacts detection performance. Covariate Shift detection is particularly critical in fields like medical imaging, where identifying distribution shifts is crucial to ensuring trustworthy predictions. Expanding such evaluations to high-stakes domains could broaden the relevance of DisCoPatch and similar models for robust, real-world applications.
While current results demonstrate the effective use of batch statistics for OOD detection, a signal-processing-focused analysis of feature propagation and suppression at each layer could provide key insights. This could reveal patterns specific to different shifts, improving model interpretability. The planned extensions include the integration of several ResNet-based architectures that use BatchNorm2D layers in the discriminator. Beyond new discriminator architectures, future work should also assess more robust reconstructors (e.g., VQ-GAN) and generators (e.g., DDPMs).
@@ 8 Conclusion
Acknowledgments
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-10021. This research was funded by the European Xecs Eureka TASTI Project.
@@ Acknowledgments
The supplementary material is organized as follows: Appendix A covers the datasets used in this work. Appendix B describes the implementation details of DisCoPatch while Appendix C covers the compute resources required for training and evaluating the models. Appendix D presents the results of the experiments on Batch normalization bias in conventional setups. Appendix E covers the impact of alternative normalization layers in the discriminator performance. Appendix F analyses in more detail the effect that multiple patch counts have on DisCoPatch’s performance. Appendix G provides detailed results on the ImageNet-1K Covariate Shift OOD detection benchmark.
@@ Appendix B Implementation Details
Table 2: Hyperparameters used for DisCoPatch’s training.
Model Lat. Dim. Hidden Dimensions lr
Full-Size 1024 32, 64, 128, 256, 512, 1024 $5e^{-5}$
Patches 1024 128, 256, 512, 1024 $8.5e^{-5}$
@@ C.2 Inference
MOODv2 (BEiTv2) 86,530,984 19.26
SCALE (ResNet-50) 25,557,032 11.27
@@
Appendix D Batch Normalization Bias
Mode Near-OOD Far-OOD
Learned Statistics 38.4/98.0 34.5/98.1
Batch Statistics BS=1 64.9/87.4 70.9/71.2
BS=16 90.2/55.5 91.1/40.0
BS=32 95.7/28.6 93.4/36.2
BS=64 99.3/2.2 96.1/23.8
BS=128 99.8/0.3 97.4/17.0
BS=256 100.0/0.0 98.1/12.5
Appendix E Alternative Normalization Layers
To better demonstrate the effect of the BatchNorm layer and batch statistics, we replaced BatchNorm with GroupNorm and InstanceNorm and retrained DisCoPatch. As shown in Table 6, GroupNorm performs the worst, while InstanceNorm shows intermediate results. These findings suggest that avoiding normalization across multiple channels is beneficial. However, BatchNorm consistently achieves the best performance, particularly under Covariate Shift, highlighting the importance of batch-level statistics.
Table 6: Comparison between different normalization layers on DisCoPatch. The results correspond to the 64-patch configuration.
Normalization Near-OOD Far-OOD Covariate Shift OOD
GroupNorm 85.0 81.0 78.5
InstanceNorm 92.8 93.7 86.2
BatchNorm 95.1+2.3 96.6+2.9 97.2+11.0
Appendix F Patch Count Imapct
Table 7 shows that the use of more patches improves the detection accuracy, reaching a plateau around 64 patches.
Table 7: OOD detection results on ImageNet-1K for multiple patch counts.
Model Near-OOD Far-OOD Covar. Shift OOD
MOODv2 88.9 97.1 70.5
SCALE 81.4 96.5 83.3
NNGuide (RegNet) 89.2 97.8 78.5
RankFeat [57] (ResNetv2-101) 89.7 94.5 91.9
ASH [7] (ResNet-50) 78.6 96.1 84.7
FDBD [29] (ResNet-50) 74.9 95.8 82.2
DisCoPatch-4 (Proposed) 90.3+0.6 92.6-5.2 93.0+1.1
DisCoPatch-16 (Proposed) 94.3+4.6 96.0-1.8 96.5+3.6
DisCoPatch-32 (Proposed) 94.8+5.1 96.4-1.4 96.9+4.0
DisCoPatch-64 (Proposed) 95.1+5.4 96.6-1.2 97.2+4.3
DisCoPatch-128 (Proposed) 95.2+5.5 96.7-1.1 97.3+4.4
DisCoPatch-512 (Proposed) 95.3+6.1 96.8-1.0 97.4+4.5
Appendix G Covariate Shift Results on ImageNet-1K